OpenStudy (anonymous):

MEDALS FOR BEST ANSWERS. Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4)?

OpenStudy (rina.r):

i think u substitute vale of x which is -3 and value of x which is 4 in this equation.

OpenStudy (anonymous):

so would it be y = -3/4x + 8. or -3/4x + 3

OpenStudy (rina.r):

no x=-3 and y=4

OpenStudy (rina.r):

4(-3)-3(4)=-12 so i have substituted values for u.

OpenStudy (anonymous):

i don't get it. those are the ones that make sense to me

OpenStudy (anonymous):

y = -3/4x + 3?

OpenStudy (rina.r):

so 4(-3)-3(4)=-12 -12-12=-12 -24=-12 so -24/-12 =2.

OpenStudy (anonymous):

so which 1?

OpenStudy (rina.r):

so that's the answer when u substitute the value in it.

OpenStudy (rina.r):

http://www.mathwarehouse.com/algebra/linear_equation/write-equation/equation-of-line-parallel-through-point.php go here it might help u understand it explain clearly so try.